# PrivacyRestore: Privacy-Preserving Inference in Large Language Models via Privacy Removal and Restoration

### v3

@@ 6.3 Empirical Privacy Protection Results
Different $d_{\chi}$-privacy Percentage for $d_{\chi}$-privacy.
We randomly select a proportion of tokens in user input to protect, denoted as the $d_{\chi}$-privacy percentage. A larger $d_{\chi}$-privacy percentage indicates a larger number of tokens being protected and a longer protected text. As illustrated in Figure 3(a) and 3(b), both prompt injection attack and attribute inference attack exhibit increased attack performance with larger $d_{\chi}$-privacy percentages, as reflected by the increase of ROUGE-L and F1 scores. These experimental results demonstrate the linear growth problem of the privacy budget in $d_{\chi}$-privacy, as proved in Section 5.1. Implementation details of attacks are presented in Appendix L.1.
Different Privacy Span Ratio $\alpha$ for PrivacyRestore.
We randomly select a proportion of privacy spans to protect, denoted as $\alpha$, where a larger $\alpha$ indicates a larger number of privacy spans being protected and longer protected text. As shown in Figure 3(c) and 3(d), the ROUGE-L scores of embedding inverse attack are stable across different $\alpha$ values on three datasets. Similarly, the F1 scores of attribute inference attack also keep stable. The length of the protected text does not impact the privacy protection capability of PrivacyRestore. The stable performance in both attack settings provides empirical evidence that our method inherently addresses the linear growth problem of privacy budget in these DP variants, as proved in Section 5.2. The trend exhibits slight fluctuations on both the Pri-NLICE and Pri-DDXPlus datasets. A detailed analysis of these fluctuations is provided in Appendix L.2.
6.4 Ablation studies
@@
In order to verify the effectiveness of Attention-aware Weighted Aggregation (AWA) component, we compare the performance and the inference efficiency between equal weighted aggregation and attention-aware weighted aggregation. Different from attention-aware weighted aggregation, equal weighted aggregation computes the meta vector by simply summing up all restoration vectors.
As shown in Table 2, the MC1, MC2, ROUGE-L, and LLM-J scores of equal weighted aggregation are all lower than those of attention-aware weighted aggregation, indicating that simply summing all restoration vectors equally degrades performance. This degradation is primarily due to the equal weights diluting the influence of critical spans while amplifying the effect of irrelevant ones. In terms of inference efficiency, the throughput difference between Attention-Aware Weighted Aggregation and Equal Weighted Aggregation is negligible. This suggests that the weight computation, as defined in Eq 4, is efficient and does not significantly impact overall throughput.
6.5 Analysis of Hyperparameter and LLM Backbone
We analyze the performance of PrivacyRestore using different numbers of edited heads $K$. In addition, we analyze the performance of PrivacyRestore using a different LLM backbone (i.e., Llama-13b-chat). Due to space limitation, we put the analysis in the Appendix I and Appendix K.
@@ 7 Conclusion
We propose PrivacyRestore which protects the privacy within user inputs during inference in online LLM inference services. PrivacyRestore achieves privacy protection by directly removing privacy spans in the user input and then restoring these privacy spans via activation steering. PrivacyRestore provides a practical and efficient solution for protecting privacy while maintaining satisfactory performance and inference efficiency. We demonstrate that PrivacyRestore inherently addresses the linear growth problem of the privacy budget found in differential privacy variants. We curate three privacy-preserving datasets covering medical and legal fields, and PrivacyRestore achieves strong performance and inference efficiency across all datasets. Additionally, we implemented two types of attacks, and the experimental results demonstrate PrivacyRestore’s robust privacy protection capabilities.
@@ Appendix A Notations
$\mathcal{S}$ All possible privacy spans.
$\mathcal{S}_{I}$ All privacy spans in user input $I$.
@@
$\mathcal{H}_{c}$ The common top-K heads set.
@@
$\mathcal{H}_{k}^{s}$ The top-K heads set of the privacy span $s$.
@@
$\mathcal{F}^{s}_{h}$ The probe of privacy span $s$ on head $h$.
$\theta^{s}_{h}$ The parameters of the probe $\mathcal{F}^{s}_{h}$.
$\sigma$ The sigmoid function.
@@
$r_{s}^{h}$ The restoration vector for privacy span $s$ on head $h$.
@@
$n$ The number of tokens in the user input.
@@
$Z_{h}$ The normalized weighted sum of restoration vectors on head $h$.
$Z_{h}$, $Z_{h}^{\prime}$ Any two normalized weighted sums.
$\mathcal{R}_{h}$ The meta vector on head $h$.
@@
$Y_{s}=\{y_{1},...,y_{m}\}$ The labels indicating whether the corresponding input contains $s$.
@@
$I=\{i_{1},...,i_{n}\}$ The tokens of the input $I$.
$i_{t}$ The $t$-th token of the user input $I$.
$\{e_{1},...,e_{n}\}$ The token embeddings of the input $I$.
$O$ The possible output sets for $I$.
$O=\{o_{1},...,o_{n}\}$ The possible output sets for tokens of the input $I$, also represented as $O$.
$\hat{I}$ The user input with all privacy spans removed.
@@
$a$ The initial output given the complete input $I$.
$\hat{a}$ The output given the incomplete input with privacy spans removed $\hat{I}$.
$Q$ , $Q^{\prime}$ Any two queries to the database.
$Q=\{q_{1},...,q_{n}\}$ The sub-queries of the query $Q$.
$G$ The possible query result sets for $Q$.
$G=\{g_{1},...,g_{n}\}$ The possible result sets for sub-queries of the query $Q$, also represented-
as $G$.
@@
$n_{ps}$ The number of tokens associated with the privacy spans in the input.
@@
$d_{e}$ The prior value of distance between token embeddings.
$d_{z}$ The prior value of distance between normalized weighted sums.
@@
Appendix B Datasets
B.1 Construction Process
We used GPT-3.5 (Ouyang et al., 2022) to classify symptoms in DDXPlus and NLICE, as well as case details in SLJA, into five levels ranging from non-sensitive to highly sensitive. The assessment prompt template is shown in Appendix O.1. A higher level indicates that the symptom or case detail is more sensitive. We define all symptoms and case details with a sensitivity level greater than 3 as privacy spans.
We assign each sample a correct answer along with three randomly selected incorrect options. For DDXPlus and NLICE, we randomly select three diagnosis results to combine with the correct diagnosis as the choices. In the SLJA dataset, we randomly select three legal judgments to pair with the correct one as the options.
B.2 Statistical Information
We show the statistics of the obtained Pri-DDXPlus, Pri-NLICE and Pri-SLJA datasets in Table 4. We tally the number of user inputs, privacy span types, and answer types. We also compute the average occurrence of privacy spans per instance. In Pri-DDXPlus and Pri-NLICE, the privacy spans are the symptoms, and the answers are the diagnoses. In Pri-SLJA, the privacy spans are the case details, and the answers are the legal judgments.
Pri-DDXPlus commonly contains more instances and more privacy symptoms types compared to Pri-NLICE and Pri-SLJA. Each sample in Pri-DDXPlus contains an average of six privacy symptoms, while samples in Pri-NLICE have an average of four privacy spans, and samples in Pri-SLJA have an average of three privacy spans.
Datasets Dataset Split User inputs Privacy Span Type Avg. Privacy Spans
Pri-DDXPlus All 7759 149 5.95
Train 5901 149 6.03
Dev 309 60 5.37
Test 1549 78 5.77
Pri-NLICE All 4062 64 4.49
Train 3282 64 4.55
Dev 130 58 4.25
Test 650 64 4.24
Pri-SLJA All 3901 142 2.67
Train 3117 142 2.56
Dev 130 95 3.21
Test 654 142 3.09
Table 4: The statistics of Pri-DDXPlus and Pri-NLICE. Average privacy symptoms indicate the average privacy spans occur in one query.
@@ Appendix C Preliminaries for Methodology
The $d_{\chi}$-privacy protection and activation steering technique are two crucial components of our method. Here, we provide an illustration of these techniques for better understanding of our method.
C.1 $d_{\chi}$-privacy protection
$d_{\chi}$-privacy (Feyisetan et al., 2019) is a variant of the differential privacy mechanism designed to protect privacy by incorporating a distance measure into the privacy budget. Its detailed definition is provided in Definition 2.1. Typically, to implement the $d_{\chi}$-privacy mechanism, noise is added to the initial embedding or vector for privacy protection, as follows:
@@ C.1 $d_{\chi}$-privacy protection
where $Z$ is the protected embedding/vector, $\mathcal{N}$ is the added noise, $\mathcal{R}$ is the protected results and $\epsilon$ is the privacy parameter of the mechanism. According to Feyisetan et al. (2019), in order to sampling the noise $\mathcal{N}$ from the distribution, we can compute as the following:
@@
where $n$ is the size of the embedding/vector and $\epsilon$ is the privacy parameter.
@@ C.2 Activation Steering Technique
Activation steering methods (Li et al., 2023c; Turner et al., 2023; Hernandez et al., 2023) control the behavior of LLM by modifying their activations during the inference stage, without incurring training costs. It serves as a crucial part of our methodology to restore privacy information during LLM inference. Typically, the attention mechanism (Vaswani et al., 2017) in LLM is responsible for capturing contextual information, and it can be expressed as:
@@
where i is the input hidden state, u is the output hidden state, $W_{q}$ is the query weight matrix, $K$ is the key of the context and $V$ is the value of the context and $d_{k}$ is the dimension of the key. Activation steering methods add some steering vectors into the output hidden state and in our methos we add the restoration meta vector into the output hidden state to restore privacy information, which can be expressed as:

### v5

@@ 6.4 Privacy Protection for Long Text
6.5 Ablation study
Attention-Aware Weighted Aggregation.
To assess the effectiveness of the Attention-Aware Weighted Aggregation (AWA) component, we compare its performance and inference efficiency against Equal Weighted Aggregation (EWA). Unlike AWA, EWA generates the meta vector by summing all restoration vectors equally. As shown in Table 3, EWA results in lower MC1, MC2, ROUGE-L, and LLM-J scores compared to AWA, indicating that equal weighting will diminish performance by amplifying irrelevant privacy spans.
@@ 6.5 Ablation study
Other Ablation Studies.
Furthermore, we evaluate PrivacyRestore’s performance with varying numbers of edited heads ($K$) and with an alternative LLM backbone (Llama-13b-chat) in Appendix U. The results in Table 15 and Figure 7 clearly demonstrate the effectiveness of our method.
6.6 Extension Analysis of PrivacyRestore
In this section, we analyze the extension PrivacyRestore to other more extreme scenarios.
Encountering Out-of-Set Privacy Spans.
Due to the long-tailed distribution of privacy spans shown in Appendix K, the core set covers most spans. We further evaluated our method when encountering out-of-set spans. Specifically, we include only a subset of privacy span types in our core set. Table 11 shows that our method still demonstrates superior performance, compared to No Restoration baseline. More implementation details are shown in Appendix S.
Users Unable to Determine Privacy Spans.
Our method follows the principle of “Information Self-Determination Rights” allowing users to determine their own privacy spans. Even when users cannot or choose not to specify these spans, our method remains effective by integrating with existing text sanitization techniques (Kan et al., 2023; Chen et al., 2023a). As shown in Table 14, our method can maintains superior performance, and details of implementation are provided in Appendix T.
@@ 7 Conclusion
We propose PrivacyRestore which protects the privacy within user inputs during inference in online LLM inference services. PrivacyRestore achieves privacy protection by directly removing privacy spans in the user input and then restoring information via activation steering. PrivacyRestore provides a practical and efficient solution for protecting privacy while maintaining satisfactory performance and inference efficiency. We demonstrate that PrivacyRestore inherently addresses the linear growth problem of the privacy budget found in $d_{\chi}$-privacy. We curate three privacy-preserving datasets covering medical and legal fields, and PrivacyRestore achieves strong performance and inference efficiency across all datasets. Additionally, we implemented various attack methods, and the attack results demonstrate PrivacyRestore’s robust privacy protection capabilities.
Limitations
This section aims to highlight the limitations of our work and provide further insights into the research in this area.
One limitation is that we only evaluate our method in the medical and legal domains and additional domains could be explored to validate its effectiveness.
Another limitation is that more attack methods could be explored to assess the privacy protection of our approach. While we have implemented most of the current advanced attack methods, to the best of our knowledge, there may be others yet to be tested. Additionally, more advanced attack methodologies may emerge in the future, which will also need to be evaluated.
Ethics Statement
We adhere to the ACL Ethics Policy and have conducted our research using publicly available repositories and datasets. In the PrivacyRestore framework, we have adhered to rigorous ethical standards to safeguard user privacy and uphold data security. All three datasets (Pri-SLJA, Pri-NLICE and Pri-DDXPlus) utilized in this research are sourced exclusively from publicly available repositories, ensuring that these datasets are devoid of any personally identifiable information and minimizing potential privacy risks. Our methodology does not access or reconstruct the original identifiable data or its sources. This ensures that the research does not infringe upon individual privacy rights.
However, due to the fact that we employed multiple LLMs in this study, such as ChatGPT, Qwen and GPT-4. The findings may be influenced by the inherent assertiveness, linguistic patterns, and diverse biases characteristic of these LLMs.
Acknowledgment
This research was supported by the National Natural Science Foundation of China (62406114,62306117,62472181), the National KeyR&D Project fromMinister of Science and Technology (2024YFA1211500), the Fundamental Research Funds for the Central Universities (2024ZYGXZR074), the Guangdong Basic and Applied Basic Research Foudation (2025A1515011413,2024A1515010220), the Guangzhou Basic and Applied Basic Research Foundation (2024A04J3681), the GJYC program of Guangzhou (2024D03J0005) and the South China University of Technology-TCL Technology Innovation Fund.
The work described in this paper was conducted in full or in part by Dr. Haoran Li, JC STEM Early Career Research Fellow, supported by The Hong Kong Jockey Club Charities Trust.
Appendix Overview
The appendix is divided into two parts: Appendix A–I provide backup theoretical explanations of PrivacyRestore, while Appendix J–X present additional experimental results on PrivacyRestore from different aspects.
@@ Appendix A Notations
$c$ A privacy span type.
$\mathcal{C}$ All possible privacy span types.
@@
$\mathcal{S}_{q}$ All privacy spans in user query $q$.
@@
$\mathcal{H}_{k}$ The common top-K heads set.
@@
$\mathcal{H}_{k}^{c}$ The top-K heads set of the privacy span type $c$.
@@
$\mathcal{F}^{c}_{h}$ The probe of privacy span type $c$ on head $h$.
$\theta^{c}_{h}$ The parameters of the probe $\mathcal{F}^{c}_{h}$.
@@
$r_{h}^{c}$ The restoration vector for privacy span type $c$ on head $h$.
@@
$n$ The number of tokens in the user query.
@@
$Z_{h}$, $Z_{h}^{\prime}$ Any two normalized weighted sums of restoration vectors on head $h$.
$\mathcal{R}$ The meta vector.
$\mathcal{R}_{h}$ The part of the meta vector for head $h$.
@@
$Y_{c}=\{y_{1},...,y_{m}\}$ The labels indicating whether the corresponding input contains the privacy span of type $c$.
@@
$\{i_{1},...,i_{n}\}$ The tokens of the input $I$.
$\{e_{1},...,e_{n}\}$ Corresponding token embeddings of the input $I$.
$O=\{o_{1},...,o_{n}\}$ The possible output sets for $I$, with each one representing a single output.
$\hat{I}$ The incomplete user input with all privacy spans removed in the training set.
@@
$a$ The output given the complete input $I$.
$\hat{a}$ The output given the incomplete input $\hat{I}$.
$q$ The user query during inference.
$\hat{q}$ The incomplete user query with all privacy spans removed during inference.
@@
$n_{ps}$ The number of tokens associated with the privacy spans in the user query.
@@
$d_{e}$ The distance between token embeddings.
$d_{z}$ The distance between normalized weighted sums.
@@
Appendix B Related Works
Here, we provide a more detailed introduction to Secure Multi-Party Computation (SMPC) and Differential Privacy (DP).
B.1 Secure Multi-Party Computation (SMPC)
Secure multi-party computation (SMPC) methods utilize multi-party encryption algorithms to enable collaborative computation among multiple parties while protecting the privacy of their data.
However, most nonlinear operations in LLMs cannot directly support secure multi-party computation.
To address this challenge, current SMPC methods focus on two optimization directions: model structure-oriented optimization and protocol-oriented optimization.
The model structure-oriented approach aims to replace SMPC-unfriendly nonlinear operations with SMPC-friendly alternatives.
For instance, MPC-Former (Li et al., 2023a) approximates nonlinear operations in Transformer using polynomials and maintains performance through model distillation.
MERGER (Liang et al., 2024) integrates previous techniques to natural language generation (NLG) tasks by bypassing embedded computation and reorganizing linear operations in Transformer modules, further enhancing computational efficiency and model performance.
In contrast, the protocol-oriented approach focuses on designing efficient SMPC operators for nonlinear operations in LLMs while preserving the original model structure.
Recent works Hao et al. (2022a); Liu and Liu (2023); Zheng et al. (2023b); Gupta et al. (2023) have improved the efficiency of nonlinear operations in privacy-preserving LLMs inference by utilizing various SMPC protocols, such as confusion circuit and function secret sharing.
Although SMPC-based methods can be applied to protect user inputs during model inference, they still suffer from large inference time overhead.
For example, inference on the RoBERTa-Base model takes 168.43 seconds (Hao et al., 2022b), making current SMPC methods impractical for online LLM inference services.
B.2 Differential Privacy (DP)
Differential Privacy (DP), as introduced by Dwork et al. (2016), is designed to protect individual privacy by preventing attackers from identifying specific participants in a dataset.
Several variants of DP have been developed to enhance privacy protection across various settings, adapting the core principles of DP to different types of data and threat models.
Notable examples include Centralized Differential Privacy (CDP), Local Differential Privacy (LDP), and $d_{\chi}$-privacy.
CDP (Dwork et al., 2016) operates under the assumption that all data has been stored in a central repository.
It guarantees that attackers cannot distinguish between any two adjacent repositories based on query results.
In contrast, LDP (Duchi et al., 2013) provides a stronger guarantee, ensuring that attackers cannot distinguish between any two adjacent inputs. Mattern et al. (2022b) and Utpala et al. (2023b) propose using paraphrasing techniques to achieve LDP on user inputs.
$d_{\chi}$-privacy (Feyisetan et al., 2019), a relaxed version of LDP, incorporates metrics that measure the similarity between inputs, allowing for more flexible control over the privacy budget. The formal definition of $d_{\chi}$-privacy Mechanism is provided in Appendix C.1.
As proposed by Mattern et al. (2022b), applying $d_{\chi}$-privacy to all tokens in user inputs, known as word-level privatization, suffers from the linear growth problem of the privacy budget.
This means that as the length of the protected text increases, the privacy protection performance of $d_{\chi}$-privacy decreases.
@@ Appendix C Preliminaries for Methodology
The $d_{\chi}$-privacy mechanism and activation steering technique are two crucial components of our method. Here, we provide a more detailed illustration of these techniques for a better understanding of our method.
C.1 $d_{\chi}$-privacy Mechanism
$d_{\chi}$-privacy mechanism (Feyisetan et al., 2019) is a variant of the differential privacy mechanism designed to protect the privacy and incorporate a distance measure into the privacy budget. The detailed definition of $d_{\chi}$-privacy is as follows:
Definition C.1.
($\bm{d_{\chi}}$-privacy mechanism). A randomized mechanism $\mathcal{M}:\mathcal{I}\to\mathcal{O}$ fulfills $\epsilon$-$d_{\chi}$-privacy if for all adjacent inputs $I,I^{\prime}\in\mathcal{I}$ and all possible outputs $O\subset\mathcal{O}$, the following condition holds:
$\mathbb{P}\left(\mathcal{M}(I)\in O\right)\leq\exp(\epsilon d_{\chi}(I,I^{\prime}))\mathbb{P}\left(\mathcal{M}(I^{\prime})\in O\right),$
where $d_{\chi}$ is a distance function defined on $\mathcal{I}$.
Numerous prior works have applied the $d_{\chi}$-privacy mechanism (Chatzikokolakis et al., 2013; Alvim et al., 2018) to word embeddings to achieve word-level privatization (Feyisetan et al., 2020, 2019; Xu et al., 2020; Bo et al., 2021). In our approach, we employ the $d_{\chi}$-privacy mechanism to protect the meta vector, preventing privacy leakage from the meta vector.
To implement the $d_{\chi}$-privacy mechanism on the meta vector/token embeddings, noise must typically be added to it, as shown below:
@@ C.1 $d_{\chi}$-privacy Mechanism
where $Z$ is the unprotected meta vector/token embeddings, $\mathcal{N}$ is the added noise, $\mathcal{R}$ is the protected meta vector/token embeddings and $\epsilon$ is the privacy parameter of the mechanism. According to Feyisetan et al. (2019), to sample the noise $\mathcal{N}$ from its distribution, we can compute it as follows:
@@
where $n$ is the size of the meta vector and $\epsilon$ is the privacy parameter.
@@ C.2 Activation Steering Technique
Activation steering methods (Li et al., 2023c; Turner et al., 2023; Hernandez et al., 2023) control the behavior of LLM by modifying their activations during the inference stage. It serves as a crucial part of our methodology to restore information contained within the removed privacy spans during LLM inference. Typically, the attention mechanism (Vaswani et al., 2017) in LLM is responsible for capturing contextual information, and it can be expressed as:
@@
where i is the input hidden state, u is the output hidden state, $W_{q}$ is the query weight matrix, $K$ is the key of the context and $V$ is the value of the context and $d_{k}$ is the dimension of the key. Activation steering methods add some steering vectors into the output hidden state and, in our methods, we add the meta vector into the output hidden state to restore information, which can be expressed as: